OpenStudy (anonymous):

Hexagon A is a regular hexagon. The total length of all the sides of the hexagon is 24 inches. Hexagon A is dilated about its center to create Hexagon B. The length of each side of Hexagon B is 4 and 4 over 5 inches. By what factor was Hexagon A dilated to create Hexagon B? 5 over 6 1 and 1 over 5 2 4 over 5

OpenStudy (cwrw238):

side of hexagon A = 24/6 = 4 inches required factor = 4 4/5 / 4 = 24/5 * 1/4 = 24/20 = 1 1/5

OpenStudy (anonymous):

thanks you

OpenStudy (cwrw238):

yw